# Projective geodesic extensions by conformal modifications in nonholonomic mechanics

### v1

@@ 1 Introduction
In mechanics, the situation changes when the problem includes nonholonomic constraints. These are non-integrable velocity-dependent restrictions of the motion of the system, such as e.g. those caused by wheels (that roll without slipping), skates (that are prevented to move in a direction perpendicular to the blade), pursuit problems, etc. Here, we consider only nonholonomic constraints that can be given in the form of linear (autonomous) equations, and that therefore can be represented by a non-integrable distribution $\mathcal{D}$ on the configuration manifold $Q$ of the mechanical system. Even though most of our results may be extended by including a potential at the appropriate places, we have chosen to limit ourselves here to purely kinetic systems, for the clarity of our exposition. If $L$ is the (kinetic energy) Lagrangian of the system, the equations of motion for such systems are the so-called Lagrange-d’Alembert equations of $(L,\mathcal{D})$. They form a set of first- and second-order ordinary differential equations whose solutions we will call the nonholonomic trajectories of $(L,\mathcal{D})$. The Lagrange-d’Alembert equations are not Euler-Lagrange equations, and are also not geodesic equations. However, it still remains useful to identify the solutions of these equations as (part of the possibly larger set of) geodesics of some metric on the configuration space. In the current paper, as in our paper [6], we investigate methods by which we can determine such a metric, which we call a geodesic extension of the nonholonomic system.
@@
In the literature, one may already find some earlier results concerning (what we now call) projective geodesic extensions. For example, in [1], the authors construct one such extension (in their Theorem 3.4), but only for the case of a Chaplygin system (a nonholonomic system with a symmetry Lie group) that is, in addition, $\varphi$-simple (in the wording of [18]). One of the main goals of this paper is to show that this very specific case is not the only one of interest: we will find (both in theory and in examples) projective geodesic extensions for nonholonomic systems that do not adhere to any of these strong assumptions.
@@
Unfortunately, the terminology Hamiltonization is over-used in the literature. In other papers such as e.g. [2, 3, 9, 15] (just to recall a few), Hamiltonization stands for the approach where one uses the symmetry group $G$ of the (Chaplygin) nonholonomic system to reduce it from a dynamical system on $Q$ to a system on the quotient manifold $Q/G$. This reduced system is then said to be Hamiltonizable if one can find a symplectic or Poisson structure on the quotient manifold for which it is Hamiltonian. To set this apart from the Hamiltonization we discussed earlier, we will call this Hamiltonian reparametrization. In the context of this type of Hamiltonization, new concepts such as $\varphi$-simplicity, dynamical gauge transformations, and invariant measures have recently been shown to play a key role [10, 18, 1, 4]. One of the main strengths of the current paper is that we are able to unify both approaches to Hamiltonization, and that we can clarify the interplay between these various subtly different notions and concepts.
For that puropose, we further explore the applicability of the conditions $(A^{\prime})$ and $(B^{\prime})$, in the event of a Chaplygin system in Section 4. Because of the symmetry properties of this type of nonholonomic systems, we can simplify these conditions and show how linear first integrals play a crucial role. The technique is then applied to the generalized nonholonomic particle. Throughout the paper, special attention is given to the subclass of geodesic extensions $\hat{g}$ which have the property that the vertical distribution $V\pi$ of the principal fibre bundle $\pi:Q\to Q/G$ is orthogonal to the constraint distribution $\mathcal{D}$. We call these type of conformal geodesic extensions $(V\pi,{\mathcal{D}})$-orthogonal. In Corollary 1 we show that, in that case, condition $(B^{\prime})$ is even redundant. This explains why condition $(B^{\prime})$ does not seem to have made any appearance yet in the literature: the assumption of $(V\pi,{\mathcal{D}})$-orthogonality is often implicitly assumed.
We have already mentioned that some of the well-known examples of Chaplygin systems possess the $\varphi$-simplicity property. In Section 5, we investigate the relation between this property and our condition $(A^{\prime})$, first at the level of the full configuration space $Q$. We use the examples to demonstrate that the class of conformal geodesic extensions is meaningful: the method of first integrals leads to conformal geodesic extensions, that are more general than those that can be obtained through the $\varphi$-simplicity of the system.
@@ 3.1 The two principal conditions
A quadratic spray (also often called affine spray) is the SODE that can be associated to a linear connection. The geodesic spray of a Riemannian metric is, in fact, the quadratic spray of its Levi-Civita connection. A quadratic spray is a special case of the more general concept of a spray (see, for example, [24] for a general reference on the geometry of sprays). Two (quadratic) sprays $\Gamma$ and $\hat{\Gamma}$ are said to be projectively equivalent if they have the same base integral curves as point sets. This means that for any base integral curve of $\tilde{\Gamma}$ there exists an orientation preserving reparametrization that transform it into a base integral curve of $\Gamma$. It is shown in e.g. [23, 24] that two sprays are projectively related if, and only if, there exists a (positively) 1-homogeneous function $P$ such that
@@
We will now make one further assumption. This assumption is, again, a generalization to the current context of that that we had also made in [6]. There, we have limited the class of the candidate metrics to those that have some special relation to the kinetic energy $g$ of the nonholonomic system: we have called a Riemannian metric $\overline{g}$ a $\mathcal{D}$-preserving modification of $g$, when
@@
A symmetric (0,2) tensor field $\hat{g}$ on $Q$ is a $\mathcal{D}$-conformal modification of a Riemannian metric $g$ on $Q$ if $\hat{g}$ is given by $e^{2F}\overline{g}$ for a function $F\in\mathcal{C}^{\infty}(Q)$ and for a Riemannian metric $\overline{g}$ which is a $\mathcal{D}$-preserving modification of $g$.
@@
In conclusion, instead of looking for a couple $(\hat{g},P)$ to obtain a projective geodesic extension, we will look for a function $F$ and a metric $\overline{g}$ such that the set of equations (7) is satisfied. It will turn out (in Proposition 1) that these conditions are independent of the choice of the components $\overline{g}_{ij}$. For this reason, we will also simply call the couple $({\overline{g}}_{ai},F)$ a projective geodesic extension (even though, technically, they provide through the freedom in $\overline{g}_{ij}$ a whole class of geodesic extensions).
@@
Let $g$ be a Riemannian metric and let $\hat{g}$ be a $\mathcal{D}$-conformal pseudo-Riemannian modification of $g$, $\hat{g}=e^{2F}\overline{g}$. Assume that ${\overline{g}}|_{\mathcal{D}^{\overline{g}}\times\mathcal{D}^{\overline{g}}}$ is non-degenerate. The conditions (9) are equivalent with the conditions
@@
    If a (0,2)-tensor field $\overline{g}$ that is a $\mathcal{D}$-preserving modification of $g$ is such that its restriction ${\overline{g}}|_{\mathcal{D}\times\mathcal{D}^{g}}=({\overline{g}}_{ai})$, together with a function $F\in\mathcal{C}^{\infty}(Q)$, satisfy the conditions $(A^{\prime})$ and $(B^{\prime})$, then there exists a projective geodesic extension of $(L,\mathcal{D})$ by $F$ and by a Riemannian metric $^{\hat{g}}$.
@@ 4.2 The generalized nonholonomic particle
for some constant $\kappa_{3}$. In fact, this is the same projective geodesic extension $(\overline{g}_{ai},F)$ that we had already encountered in Section 3.2, under the identification $\kappa_{1}=a_{2}$ and $\kappa_{3}=-a_{1}$.
@@
from which $K=\kappa_{4}\rho^{\prime}$ (with $\kappa_{4}$ a constant). We conclude that, for any solution of the SODE in $\rho$, we can find the following corresponding conformal factor $F$:
@@ 5.1 Relation to projective geodesic extensions on the configuration manifold
    (that we had already encountered in the previous section), are the ideal candidates. So far, this natural assumption remained somewhat hidden in the literature [1] on this topic.
@@
The conclusion is that $\varphi$-simplicity induces, in a very specific way, a projective geodesic extension. It also explains why condition $(B^{\prime})$ does not seem to have made any appearance yet in the literature: the assumption of $(V\pi,{\mathcal{D}})$-orthogonality is often implicitly assumed. The result in Proposition 3 encompasses Theorem 3.4 of [1]. It is even a bit more general, in the sense that [1] only describes one specific projective geodesic extension metric, by fixing an appropriate ${\overline{g}}_{ij}$, while we know that this is not necessary (in view of Propostion 1).
@@ 5.3 The two-wheeled carriage
From this, we get that the two-wheeled carriage is indeed $\varphi$-simple, with
@@
Together with what we had said about this example in Section 3.3, we now reach the following conclusions:
  • For any $\ell$, the couple $(\overline{g}_{ai}=-G_{ai},F=\varphi\circ\pi)$ gives a projective geodesic extension of the two-wheeled carriage (according to Proposition 3).
  • When $\ell$ has the special value $\ell_{1}\neq 0$ we have actually found two distinct projective geodesic extensions: $(\overline{g}_{ai}=-G_{ai},F=\varphi\circ\pi)$ and $(\overline{g}_{ai}=\hat{g}_{ai},F=0)$ with $\hat{g}_{ai}$ as in expressions (11). The second is unrelated to the $\varphi$-simplicity property of this Chaplygin system.
  • When $\ell=0$, the expressions for $G_{ai}$ overlap with those in (11), taking the extra minus-sign into account. In that case also $\varphi=0$, and this is in agreement with $F=0$. We have therefore found only one projective geodesic extension.
@@ 5.4 Reduction of the conditions for a projective geodesic extension
Suppose that $G$ is a connected Lie group. A tensor field $t$ is $G$-invariant if and only if $\mathcal{L}_{V_{i}}t=0$, for any infentisesimal generator $V_{i}$.
@@ 5.5 Relation to projective geodesic extensions on the reduced manifold
We can now use these tensor fields to give an equivalent condition for $\varphi$-simplicity. The condition (ii) below appears already in some way in [10, 22]. With Proposition 5, we prove that it is in fact, nothing more than the $\varphi$-simplicity property in hidden form.
@@
    $\varphi$ satisfies the $\varphi$-simplicity property, $\quad(\Leftrightarrow\quad d\varphi\wedge d\theta_{l}=\Xi^{G})$.
@@
    $\begin{cases}\textrm{The function $f=\varphi$ satisfies }(A^{\prime})^{G}\quad(\Leftrightarrow\quad d\varphi\wedge d\theta_{l}=\gamma^{G}),\\[5.69054pt] \displaystyle\sum_{cycl:\tilde{X},\tilde{Y},\tilde{Z}}g^{red}(\tilde{X},{\mathcal{T}}(\tilde{Y},\tilde{Z}))=0,\,\,\forall\tilde{X},\tilde{Y},\tilde{Z}\in\mathcal{X}(Q/G).\end{cases}$
@@ 6.1 Sufficient conditions
From Lemma 6, we know that condition $(A^{\prime})$ is equivalent with $df\wedge\theta_{l}=\gamma$. We need to show that $e^{f}(\omega_{l}-\gamma)$ is closed. This is the case:
@@
Given that also $\iota_{e^{-f}\Gamma^{red}}(e^{f}\tilde{\tilde{\omega}}^{G})=dE_{l}$, we see that $e^{-f}\Gamma^{red}$ is a Hamiltonian vector field with respect to the closed $e^{f}\tilde{\tilde{\omega}}^{G}$. We can now compute the Lie derivative of $e^{f}\tilde{\tilde{\omega}}^{G}$ with respect to this vector field $e^{-f}\Gamma^{red}$:
@@ 6.2 Equivalent condition
Let $\beta$ be the contraction (23) of the Chaplygin system. The following statements are locally equivalent:
@@
    There exists a $f$ that satisfies $(A^{\prime})^{G}$ (i.e. there exist a projective geodesic extension $(\overline{g}_{ai}=-G_{ai},F=f\circ\pi)$).
@@
Let $\beta$ be the contraction (23) of the Chaplygin system. The following statements are locally equivalent:
@@
    There exists a $\varphi\in{\mathcal{C}}^{\infty}(Q/G)$ for which the system is $\varphi$-simple.
@@
When we work in the case of a $(V\pi,{\mathcal{D}})$-orthogonal projective geodesic extension, we have reached the following conclusions:
Proposition 12.
Suppose that $\overline{g}_{ai}=-G_{ai}$. Then the following (local) statements hold.
@@
When the dimension of $Q/G$ is 2, there are no 3-forms on $Q/G$, and therefore the cyclic condition is void and $\Theta^{G}=0$. Also, $\beta=-R^{2}_{12}dq^{1}+R^{1}_{12}dq^{2}$, $\theta_{l}=(g_{11}{\dot{q}}^{1}+g_{12}{\dot{q}}^{2})dq^{1}+(g_{12}{\dot{q}}^{1}+g_{22}{\dot{q}}^{2})dq^{2}$, $\gamma^{G}=\Xi^{G}=-(R^{1}_{12}((g_{11}{\dot{q}}^{1}+g_{12}{\dot{q}}^{2}))+R^{2}_{12}(g_{12}{\dot{q}}^{1}+g_{22}{\dot{q}}^{2}))dq^{1}\wedge dq^{2}$, and therefore automatically $\beta\wedge\theta_{k}\equiv\gamma^{G}$. So, the only condition that remains is $d\beta=0$. ∎

### v2

@@ 1 Introduction
In mechanics, the situation changes when the problem includes nonholonomic constraints. These are non-integrable velocity-dependent restrictions of the motion of the system, such as e.g. those caused by wheels (that roll without slipping), skates (that are prevented to move in a direction perpendicular to the blade), pursuit problems, etc. Here, we consider only nonholonomic constraints that can be given in the form of linear (autonomous) equations, and that therefore can be represented by a non-integrable distribution $\mathcal{D}$ on the configuration manifold $Q$ of the mechanical system. Even though most of our results may be extended by including a potential at the appropriate places, we have chosen to limit ourselves here to purely kinetic systems, for the clarity of our exposition. If $L$ is the (kinetic energy) Lagrangian of the system, the equations of motion for such systems are the so-called Lagrange-d’Alembert equations of $(L,\mathcal{D})$. They form a set of first- and second-order ordinary differential equations whose solutions we will call the nonholonomic trajectories of $(L,\mathcal{D})$. The Lagrange-d’Alembert equations are not Euler-Lagrange equations, and are also not geodesic equations. However, it still remains useful to identify the solutions of these equations as (part of the possibly larger set of) geodesics of some metric on the configuration space. In the current paper, as in our paper [6], we investigate methods by which we can determine such a metric. If that can be achieved, we will speak of a ‘geodesic extension’ of the nonholonomic dynamics (even though this question seems to be dubbed ‘the kinetic Lagrangianization problem’ in the conclusions of the paper [1]).
@@
In the literature, one may already find some earlier results concerning (what we now call) projective geodesic extensions. For example, in [1], the authors construct one such extension (in their Theorem 3.4), but only for the case of a Chaplygin system (a nonholonomic system with a symmetry Lie group) that is, in addition, $\varphi$-simple (in the wording of [19]). In this paper we start from scratch, at the level of an arbitrary kinetic nonholonomic system, and we use a different approach. One of the main goals of this paper is to show that the very specific case of Theorem 3.4 in [1] is not the only one of interest: we will find (both in theory and in examples) projective geodesic extensions for nonholonomic systems that do not adhere to any of these strong assumptions.
@@
Unfortunately, the terminology Hamiltonization is over-used in the literature. In other papers such as e.g. [2, 3, 9, 15] (just to recall a few), Hamiltonization stands for the approach where one uses the symmetry group $G$ of the (Chaplygin) nonholonomic system to reduce it from a dynamical system on $Q$ to a system on the quotient manifold $Q/G$. This reduced system is then said to be Hamiltonizable if one can find a symplectic or Poisson structure on the quotient manifold for which it is Hamiltonian. To set this apart from the Hamiltonization we discussed earlier, we will call this Hamiltonian reparametrization. In the context of this type of Hamiltonization, new concepts such as $\varphi$-simplicity, dynamical gauge transformations, and invariant measures have recently been shown to play a key role [1, 4, 10, 19, 16, 28]. One of the main strengths of the current paper is that we are able to unify both approaches to Hamiltonization, and that we can clarify the interplay between these various subtly different notions and concepts.
For that puropose, we further explore the applicability of the conditions $(A^{\prime})$ and $(B^{\prime})$, in the event of a Chaplygin system in Section 4. Because of the symmetry properties of this type of nonholonomic systems, we can simplify these conditions (see Prop 2) and at the end of the section we show how linear first integrals can be employed. The technique is then applied to the generalized nonholonomic particle. Throughout the paper, special attention is given to the subclass of geodesic extensions $\hat{g}$ which have the property that the vertical distribution $V\pi$ of the principal fibre bundle $\pi:Q\to Q/G$ is orthogonal to the constraint distribution $\mathcal{D}$. We call these type of conformal geodesic extensions $(V\pi,{\mathcal{D}})$-orthogonal. The above mentioned result of [1] falls into this category. In Corollary 1 we show that, in that case, condition $(B^{\prime})$ is even redundant. This explains why condition $(B^{\prime})$ does not seem to have made any appearance yet in the literature: the assumption of $(V\pi,{\mathcal{D}})$-orthogonality is often implicitly assumed.
We have already mentioned that some of the well-known examples of Chaplygin systems possess the $\varphi$-simplicity property. In Section 5, we investigate the relation between this property and our condition $(A^{\prime})$, first at the level of the full configuration space $Q$. We use the examples to demonstrate that the class of conformal geodesic extensions is meaningful: The metrics we had found in Section 3.2 lead to conformal geodesic extensions, that are more general than those that can be obtained through the $\varphi$-simplicity of the system.
@@
Throughout this paper, the concepts we introduce are globally defined. However, in specific examples one often finds only local expressions. In Section 5.3 we discuss some consequences of this in some detail.
@@ 3.1 The two principal conditions
A quadratic spray (also often called affine spray) is the SODE that can be associated to a linear connection. The geodesic spray of a Riemannian metric is, in fact, the quadratic spray of its Levi-Civita connection. A quadratic spray is a special case of the more general concept of a spray (see, for example, [26] for a general reference on the geometry of sprays). Two (quadratic) sprays $\Gamma$ and $\hat{\Gamma}$ are said to be projectively equivalent if they have the same base integral curves as point sets. This means that for any base integral curve of $\hat{\Gamma}$ there exists an orientation preserving reparametrization that transform it into a base integral curve of $\Gamma$. It is shown in e.g. [25, 26] that two sprays are projectively related if, and only if, there exists a (positively) 1-homogeneous function $P$ such that
@@
We will now make one further assumption. This assumption is, again, a generalization to the current context of that that we had also made in [6]. There, we have limited the class of the candidate metrics to those that have some special relation to the kinetic energy $g$ of the nonholonomic system: we have called a symmetric $(0,2)$-tensor field $\overline{g}$ a $\mathcal{D}$-preserving modification of $g$, when
@@
A symmetric (0,2) tensor field $\hat{g}$ on $Q$ is a $\mathcal{D}$-conformal modification of a Riemannian metric $g$ on $Q$ if $\hat{g}$ is given by $e^{2F}\overline{g}$ for a function $F\in\mathcal{C}^{\infty}(Q)$ and for a symmetric $(0,2)$-tensorfield $\overline{g}$ which is a $\mathcal{D}$-preserving modification of $g$.
@@
In conclusion, instead of looking for a couple $(\hat{g},P)$ to obtain a projective geodesic extension, we will look for a function $F$ and a Riemannian metric $\overline{g}$ such that the set of equations (7) is satisfied. It will turn out (in Proposition 1) that these conditions are independent of the choice of the components $\overline{g}_{ij}$. For this reason, we will also simply call the couple $({\overline{g}}_{ai},F)$ a projective geodesic extension (even though, technically, they provide through the freedom in $\overline{g}_{ij}$ a whole class of geodesic extensions).
@@
Let $g$ be a Riemannian metric and let $\hat{g}$ be a pseudo-Riemannian metric which is a $\mathcal{D}$-conformal modification of $g$, $\hat{g}=e^{2F}\overline{g}$. Assume that ${\overline{g}}|_{\mathcal{D}^{\overline{g}}\times\mathcal{D}^{\overline{g}}}$ is non-degenerate, where $\mathcal{D}^{\overline{g}}$ stands for the orthogonal of $\mathcal{D}$ w.r.t. $\overline{g}$. The conditions (9) are equivalent with the conditions
@@
    If a pseudo-Riemannian metric $\overline{g}$ that is a $\mathcal{D}$-preserving modification of $g$ is such that its restriction ${\overline{g}}|_{\mathcal{D}\times\mathcal{D}^{g}}=({\overline{g}}_{ai})$, together with a function $F\in\mathcal{C}^{\infty}(Q)$, satisfy the conditions $(A^{\prime})$ and $(B^{\prime})$, then there exists a projective geodesic extension of $(L,\mathcal{D})$ by $F$ and by a Riemannian metric $^{\hat{g}}$.
@@ 4.2 The generalized nonholonomic particle
for some constant $\kappa_{3}$ (where $\kappa_{1}$ and $\kappa_{3}$ are assumed not both to be zero). In fact, this is the same projective geodesic extension $(\overline{g}_{ai},F)$ that we had already encountered in Section 3.2, under the identification $\kappa_{1}=a_{2}$ and $\kappa_{3}=-a_{1}$.
@@
from which $K=\kappa_{4}\rho^{\prime}$ (with $\kappa_{4}$ a constant). We conclude that, for any solution of the SODE in $\rho$, we can find the following local conformal factor $F$:
@@ 5.1 Relation to projective geodesic extensions on the configuration manifold
    (that we had already encountered in the previous section) are the ideal candidates. This natural assumption also lies at the basis of the specific choice that has been made in Theorem 3.4 in [1].
@@
The conclusion is that $\varphi$-simplicity induces, in a very specific way, a projective geodesic extension. It also explains why condition $(B^{\prime})$ does not seem to have made any appearance yet in the literature: the assumption of $(V\pi,{\mathcal{D}})$-orthogonality is often implicitly assumed. The result in Proposition 3 encompasses Theorem 3.4 of [1]. In their Remark 3.5 the authors of [1] had already mentioned that their choice is not unique. Now, we know that the remaining freedom is encoded in condition $(B^{\prime})$.
@@ 5.3 The two-wheeled carriage
From this, we get that the two-wheeled carriage is $\varphi$-simple, with
@@
Unfortunately, the function $\varphi$ is only locally defined since it is not periodic in $\psi_{1}$ and $\psi_{2}$. Together with what we had said about this example in Section 3.3, we now reach the following conclusions:
  • For any $\ell$, the couple $(\overline{g}_{ai}=-G_{ai},F=\varphi\circ\pi)$ gives a local projective geodesic extension of the two-wheeled carriage (according to Proposition 3).
  • When $\ell$ has the special value $\ell_{1}\neq 0$ we have actually found two distinct projective geodesic extensions: $(\overline{g}_{ai}=-G_{ai},F=\varphi\circ\pi)$ and $(\overline{g}_{ai}=\hat{g}_{ai},F=0)$ with $\hat{g}_{ai}$ as in expressions (11). The second is global and unrelated to the $\varphi$-simplicity property of this Chaplygin system.
  • When $\ell=0$, the expressions for $G_{ai}$ overlap with those in (11), taking the extra minus-sign into account. In that case also $\varphi=0$, and this is in agreement with $F=0$. We have therefore found only one global projective geodesic extension.
@@ 5.4 Reduction of the conditions for a projective geodesic extension
Suppose that $G$ is a connected Lie group. A tensor field $t$ is $G$-invariant if and only if $\mathcal{L}_{V_{i}}t=0$, for any infinitesimal generator $V_{i}$.
@@ 5.5 Relation to projective geodesic extensions on the reduced manifold
We can now use these tensor fields to give an equivalent condition for $\varphi$-simplicity. The condition (ii) below appears already in some way in [10, 19, 24]. With Proposition 5, we prove that it is in fact, nothing more than the $\varphi$-simplicity property in hidden form.
@@
    $\varphi$ satisfies the $\varphi$-simplicity property, $\quad(\Leftrightarrow\quad d\varphi\wedge\theta_{l}=\Xi^{G})$.
@@
    $\begin{cases}\textrm{The function $f=\varphi$ satisfies }(A^{\prime})^{G}\quad(\Leftrightarrow\quad d\varphi\wedge\theta_{l}=\gamma^{G}),\\[5.69054pt] \displaystyle\sum_{cycl:\tilde{X},\tilde{Y},\tilde{Z}}g^{red}(\tilde{X},{\mathcal{T}}(\tilde{Y},\tilde{Z}))=0,\,\,\forall\tilde{X},\tilde{Y},\tilde{Z}\in\mathcal{X}(Q/G).\end{cases}$
@@ 6.1 Sufficient conditions
From Corollary 2, we know that condition $(A^{\prime})$ is equivalent with $df\wedge\theta_{l}=\gamma$. We need to show that $e^{f}(\omega_{l}-\gamma)$ is closed. This is the case:
@@
Given that also $\iota_{e^{-f}\Gamma^{red}}(e^{f}\tilde{\tilde{\omega}}^{G})=dE_{l}$, we see that $e^{-f}\Gamma^{red}$ is a Hamiltonian vector field with respect to the closed form $e^{f}\tilde{\tilde{\omega}}^{G}$. We can now compute the Lie derivative of $e^{f}\tilde{\tilde{\omega}}^{G}$ with respect to this vector field $e^{-f}\Gamma^{red}$:
@@ 6.2 Equivalent condition
Let $\beta$ be the contraction (23) of the Chaplygin system. The following statements are equivalent:
@@
    There exists a locally defined $f$ that satisfies $(A^{\prime})^{G}$ (i.e. there exist a projective geodesic extension $(\overline{g}_{ai}=-G_{ai},F=f\circ\pi)$).
@@
Let $\beta$ be the contraction (23) of the Chaplygin system. The following statements are equivalent:
@@
    There exists a locally defined $\varphi\in{\mathcal{C}}^{\infty}(Q/G)$ for which the system is $\varphi$-simple.
@@
The previous and the next statements refer to locally defined functions because we make use of the Lemma of Poincare. In the global case only one implication holds.
Theorem 1.
Suppose that $\overline{g}$ is $(V\pi,{\mathcal{D}})$-orthogonal. Then the following (local) statements hold.
@@
When the dimension of $Q/G$ is 2, there are no 3-forms on $Q/G$, and therefore the cyclic condition is void and $\Theta^{G}=0$. Also, $\beta=-R^{2}_{12}dq^{1}+R^{1}_{12}dq^{2}$, $\theta_{l}=(g_{11}{\dot{q}}^{1}+g_{12}{\dot{q}}^{2})dq^{1}+(g_{12}{\dot{q}}^{1}+g_{22}{\dot{q}}^{2})dq^{2}$, $\gamma^{G}=\Xi^{G}=-(R^{1}_{12}((g_{11}{\dot{q}}^{1}+g_{12}{\dot{q}}^{2}))+R^{2}_{12}(g_{12}{\dot{q}}^{1}+g_{22}{\dot{q}}^{2}))dq^{1}\wedge dq^{2}$, and therefore automatically $\beta\wedge\theta_{l}\equiv\gamma^{G}$. So, the only condition that remains is $d\beta=0$. ∎